Write an equation of the line perpendicular to the given line and with the given y-intercept.
y=-3x+2
b=4

Answer by jim_thompson5910(35256)

The slope of y=-3x+2 is -3 = -3/1

Flip the fraction: -3/1 ---> -1/3

Flip the sign: -1/3 ---> +1/3 = 1/3

So the slope of this perpendicular line is 1/3

m = 1/3
b = 4

Plug those values in to get

y = mx + b

y = (1/3)x + 4